Please help me with this problem. Evaluate (-6y) to the second power and -6y to the second power when y=2. I would really appreciate some help.
Thank you,

Answer by Alan3354(30993)

Please help me with this problem. Evaluate (-6y) to the second power and -6y to the second power when y=2.
---------------------
(-6y)^2 = 36y^2 = 144
-6y^2 = -6*4 = -24